# Odd-parity Magnetism Driven by Antiferromagnetic Exchange

### v1

@@ Abstract
Realizing odd-parity, time-reversal-preserving, non-relativistic spin splitting is a central goal for spintronics applications. We propose a group-theory-based microscopic framework to induce odd-parity spin splitting from coplanar antiferromagnetic (AFM) states without spin-orbit coupling (SOC). We develop phenomenological models for 421 conventional period-doubling AFM systems in non-symmorphic space groups and construct minimal microscopic models for 119 of these. We find that these AFM states can attain three possible competing ground states. These ground states all break symmetries in addition to those broken by the usual AFM order. Specifically, they give rise to either odd-parity spin-splitting, nematic order, or scalar odd-parity order related to multiferroicity. Our microscopic theories reveal that the odd-parity spin-splitting energy scale is generically large and further reveal that the scalar odd-parity order gives a non-zero Berry curvature dipole without SOC. We identify 67 materials in the Magndata database for which our theory applies. We provide DFT calculations on FeSe that reveal an $h$-wave spin splitting consistent with our symmetry arguments and apply our microscopic model to determine the non-relativistic Edelstein response for CeNiAsO.
@@
The odd-parity spin splitting is induced by ${\bf S_{1}}\times{\bf S_{2}}$. This induced order is time-reversal invariant, as time-reversal symmetry flips both spins, and is translationally invariant. The inversion symmetry acts non-trivially on ${\bf S_{1}}$ and ${\bf S_{2}}$ (into blue dashed arrows). In the class 1, inversion interchanges ${\bf S_{1}}$ and ${\bf S_{2}}$. In the class 2, inversion leaves ${\bf S_{1}}$ unchanged but flips ${\bf S_{2}}$. In both cases, the resulting spin splitting ${\bf S_{1}}\times{\bf S_{2}}$ is odd under inversion.
@@
Figure 2: (Left) FeSe coplanar magnetic state (Center) DFT band dispersion along the momentum cut ±$\pm(\pi,\pi/2,\pi)$. (Right) Fermi surface ($k_{z}=0.4\pi$ section) exhibiting $h$-wave spin splitting.
To test our symmetry-based arguments, we have examined if the predicted $h$-wave $k_{x}k_{y}k_{z}(k_{x}^{2}-k_{y}^{2})\sigma_{z}$ spin-splitting appears within a DFT calculation for FeSe with coplanar order at the M-point $(\pi,\pi,0)$. Collinear AFM ordering at the M point is known to generate the nematic order observed in many Fe-based materials Kreisel et al. (2020). A coplanar magnetic order is sometimes observed in this class of materials Ding et al. (2017), however, this coplanar state appears in Fe-based materials with space group 139, not space group 129 as considered here. Nevertheless, this observation opens the possibility that such a coplanar state may be stabilized in a Fe-based material with space group 129. As shown in Fig. 2, the DFT results for this coplanar state agree with our symmetry prediction. Furthermore, this calculation reveals that the spin-splitting, even for a highly nodal $h$-wave state, can be as large as 0.1 eV.
Phenomenological model: We develop a Landau theory to describe the 421 different AFM phase transitions and analyze all the possible induced orders. In the absence of SOC, the free energy density is
@@
The derivation is in the SM. $\epsilon_{abc}$ is the antisymmetric tensor. The nontrivial k-odd dependence from $t_{x}t_{y}$ carries the site symmetry, as the AFM state distinguishes the two sublattices (state 1 in Fig. 3). This k-odd Berry curvature can generate non-zero Berry curvature dipoles $\langle d\Omega^{ij}/dk_{m}\rangle_{FS}$, which drives non-relativistic nonlinear I-V transport coefficients Sodemann and Fu (2015); Kaplan et al. (2024). In addition, this state will also exhibit ferroelectric-like orderCheong and Mostovoy (2007).

### v2

@@ Abstract
Realizing odd-parity, time-reversal-preserving, non-relativistic spin splitting is a central goal for spintronics applications. We propose a group-theory-based microscopic framework to induce odd-parity spin splitting from coplanar antiferromagnetic (AFM) states without spin-orbit coupling (SOC). We develop phenomenological models for 421 conventional period-doubling AFM systems in non-symmorphic space groups and construct minimal microscopic models for 119 of these. We find that these AFM states can attain three possible competing ground states. These ground states all break symmetries in addition to those broken by the usual AFM order. Specifically, they give rise to either odd-parity spin-splitting, nematic order, or scalar odd-parity order related to multiferroicity. Our microscopic theories reveal that the odd-parity spin-splitting energy scale is generically large and further reveal that the scalar odd-parity order gives a non-zero Berry curvature dipole without SOC. We identify 67 materials in the Magndata database for which our theory applies. We provide DFT calculations on Fe-based materials that reveal an $h$-wave spin splitting consistent with our symmetry arguments and apply our microscopic model to determine the non-relativistic Edelstein response for CeNiAsO.
@@
Here ${\bf S_{1}}\times{\bf S_{2}}$ is an important quantity, as we discuss later, it generates the odd-parity spin-splitting. This induced order is time-reversal invariant, as time-reversal symmetry flips both spins, and is translationally invariant. The inversion symmetry acts non-trivially on ${\bf S_{1}}$ and ${\bf S_{2}}$ (into blue dashed arrows). In the class 1, inversion interchanges ${\bf S_{1}}$ and ${\bf S_{2}}$. In the class 2, inversion leaves ${\bf S_{1}}$ unchanged but flips ${\bf S_{2}}$. In both cases, the resulting spin splitting ${\bf S_{1}}\times{\bf S_{2}}$ is odd under inversion. Here ${\bf S_{1}}\times{\bf S_{2}}$ can also be interpreted as a non-relativistic version of the Dzyaloshinsky-Moriya vector Dzyaloshinsky (1958); Moriya (1960).
@@
Figure 2: (Left) FeSe coplanar magnetic state (Center) DFT band dispersion along the momentum cut ±$\pm(\pi,0.4\pi,0.5\pi)$. (Right) Fermi surface ($k_{z}=0.4\pi$ section) exhibiting $h$-wave spin splitting. The band spin induced is represented by a red-white-blue color code.
To test our symmetry-based arguments, we have examined if the predicted $h$-wave $k_{x}k_{y}k_{z}(k_{x}^{2}-k_{y}^{2})\sigma_{z}$ spin-splitting appears within a DFT calculation for FeSe with coplanar order at the M-point $(\pi,\pi,0)$. Collinear AFM ordering at the M point is known to generate the nematic order observed in many Fe-based materials Kreisel et al. (2020). A coplanar magnetic order is sometimes observed in this class of materials Ding et al. (2017); Stadel et al. (2022). In particular, this coplanar state appears in LaFeAs1-xPx Stadel et al. (2022) with space group 129, providing a materials realization of this state. As shown in Fig. 2, the DFT results for this coplanar state agree with our symmetry prediction. Furthermore, this calculation reveals that the spin-splitting, even for a highly nodal $h$-wave state, can be as large as 0.1 eV.
Phenomenological model: We apply Landau theory to describe the 421 different AFM phase transitions and analyze all the possible induced orders. In the absence of SOC, the free energy density of the 2D IR is Fernandes et al. (2016)
@@
The derivation is in the SM. $\epsilon_{abc}$ is the antisymmetric tensor. The nontrivial k-odd dependence from $t_{x}t_{y}$ carries the site symmetry, as the AFM state distinguishes the two sublattices (state 1 in Fig. 3). This k-odd Berry curvature can generate non-zero Berry curvature dipoles $\langle d\Omega^{ij}/dk_{m}\rangle_{FS}$, which drives non-relativistic nonlinear I-V transport coefficients Sodemann and Fu (2015); Kaplan et al. (2024). In addition, this state is related to multiferroicity Cheong and Mostovoy (2007); Matsuda et al. (2024) and will also exhibit ferroelectric-like orderCheong and Mostovoy (2007).